Hi,
My problem is about inequality
Abu ran on-third of an hour and then walked for half an hour at the rate of 3 km/h. In all, he had covered 4 km. how fast did he run?
PLSSSS. It's an assignment which is not yet corrected

Answer by stanbon(75887)

Abu ran on-third of an hour and then walked for half an hour at the rate of 3 km/h. In all, he had covered 4 km. how fast did he run?
---------------
Running DATA:
time = (1/3) hr ; rate = x km/h ; distance = rt = 3x km
--------------
Walking DATA:
time = (1/2) hr ; rate = 3 km/h ; distance = rt = 3(1/2) = 3/2 km
---------------------------
EQUATION:
distance + distance = 4 km
3x + (3/2) = 4
3x = 5/2
x = 5/6 km/h (rate at which he ran)
=======================================
Cheers,
Stan H.